Remember ...

Overall Goal

Exploit Web technologies for course-grain distributed computing by integrating them with a careful mix of interpreted and compiled environments for HPC programming languages.

WebHPF & PCRC Relationship

WebHPF = PCRC + an attractive front-end + low latency Web compatible message passing systems

Kivanc Dincer November 16, 1996

PCRC.

A Three Layer Architecture for WebHPL

• Layer 1: – Define computation and communication primitives, based on Web standards.

– Construct a World-Wide Virtual Machine(WWVM) for heterogeneous and

distributed high performance computing.

• Layer 2: – more advanced client paradigms

– interchange of different data formats and script languages.

– HPFCL - for coarse grained control-parallel HPF modules.

• Layer 3: – User level programming and problem solving environments(PSE)

– Example PSEs: PCRC, Global Arrays

Kivanc Dincer November 16, 1996 WebHPL 4

Tasks Schedule

Kivanc Dincer November 16, 1996

Current PCRC Tasks:

• Boston – NPAC software integration.

• A multi-language PCRC runtime library: – A combination of HPF/F90D Fortran and C runtime libraries.

– incremental updates towards a more advanced PCRC library in C.